Fermi National Laboratory


Accelerator Update

Wednesday June 23
The day shift began with Operations monitoring store 3588, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

NuMI experts began a mutibatch study at 9 AM.

At 10:03 AM, the Tevatron (TeV) suffered a sector E11 quench due to a correction magnet (HF48) power supply being put into local, which caused an orbit distortion.

Operators made the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls ready for access at 10:22 AM.

MCenter had turn off earlier, with beam sent to the Switchyard (SY) dump. Operators disabled all SY beam at 11:13 AM.

Both CDF and D0 came out of access at 1 PM and Operations began the TeV turn on.

At 1:20 PM, Fermi suffered from a ComEd power glitch.

At 1:22 PM, Operations reported that Linac was down as well as Booster RF due to the glitch. Both Pbar and Main Injector (MI) were still up and ramping. The TeV was turned off, but it hadn't tripped.

Technicians brought both Linac and Booster up so that could Pbar resumed stacking at 1:37 PM.

Operations put the TeV through a dry squeeze at 2:05 PM and into shot setup at 2:33 PM.

At 3:09 PM, Operations reported having beam problems with MI. Expert investigated.

A Preaccelerator (Preacc) expert began tuning on the H- Source at 3:46 PM. The intensity had dropped a little.

Operations established store 3589 at 5:06 PM with an initial luminosity of 82.3E30.

Pbar resumed stacking around 5:30 PM.

Operations began trying to send beam to the Meson MTest experiment at 6:40 PM.

At 7:42 PM, a Booster RF station (BRF12) tripped off for the first of three times in eight minutes. It reset okay.

Operations began sending beam to MCenter at 11:18 PM.

At 11:20 PM, BRF12 tripped off twice within a minute. It reset okay.

Thursday June 24
The midnight shift began with Operations monitoring store 3589, with Pbar stacking, with Recycler experts conducting studies, and with MiniBooNE, MCenter, and MTest taking beam.

Recycler experts finished their studies at 2:17 AM.

At 4:56 AM BRF12 tripped off. Operations reported that this was the third time it had tripped off this shift. It reset okay.

Operations turned off the Linac at 7:26 AM in preparation for the planned maintenance access today. Booster was off by 7:37 AM, as well as MI.

Linac techs completed their work and came out of access. Operations began turning the machine back on around 9 AM.

Booster accesses finished at 10:52 AM.

At 11:41 AM, Operations reported problems with the Linac beam due to a RF station (LRF2). Experts investigated and said nothing was wrong with LRF2.

A Booster RF station (BRF11) tripped off 12:36 PM and then again ten minutes later. An expert investigated.

At 12:45 PM, Operations reported that they were still having problems with the Linac beam. Experts continued their investigation and adjusted the voltage for a quadrupole power supply (Q4) at 1 PM.

MI work finished and Pbar resumed stacking at 1:16 PM.

Operations reestablished beam to MiniBooNE at 3:09 PM and to SY a short time later.

Operations established beam to MCenter at 4:24 PM.

At 6:22 PM, BRF12 tripped off for the first of three times in four minutes. It reset okay.

TeV experts began an end-of-store study at 9 PM.

BRF12 tripped off at 10:02 PM. It reset okay.

Operations terminated store 3589 at 10:17 PM.

BRF12 tripped off at 11:11 PM. It reset okay.

Friday June 25
The midnight shift began with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE, MTest, and MCenter taking beam.

At 12:26 AM, Operations reported that BRF12 had tripped off twice.

Operations prepared D0 for a half hour access at 1 AM.

Operations put the TeV through a dry squeeze at 1:50 AM and into shot setup at 2:12 AM.

Operations established store 3592 at 5 AM with an initial luminosity of 83.29E30. Pbar resumed stacking and Operations resumed sending beam to Meson experiments.

At 8 AM, Operations reported that they had had too many BRF12 tripped to either count or log.

Plans for Today and this Weekend
The plans for today and this weekend are to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/25/2004   email Fermilab